OpenStudy (anonymous):

Assume the random variable X is normally distributed with mean u = 50 and standard deviation o = 7. Compute the probability. Be sure to draw a normal curve with the area corresponding to the probability shaded. P(X > 43)

OpenStudy (anonymous):

I was able to recognize the shaded area but I'm not sure how to find the probability. I know to go to 2nd vars, but that's it

OpenStudy (anonymous):

damn rly need help on this; ill buy pizza and beer for anyone who wants 2 answer this and turn tf up

OpenStudy (anonymous):

In this case \(x=43\). We want to change it to \(z\) score. We use \[ z=\frac{x-\mu}{\sigma} = \frac{53-50}{7} = -\frac{3}{7} \]

OpenStudy (anonymous):

Just look up \(-3/7\) in a \(z\) table and you are done. http://www.wolframalpha.com/input/?i=z+score+-3%2F7 Use right-tailed p-value for probability of being greater.